OpenStudy (anonymous):

Triangle JKL where angle K is a right angle. KL measures 34; JL measures x; angle L measures 27 degrees. Answer 30.29 38.16 66.73 74.89

OpenStudy (anonymous):

38.16 because \[ \cos(27) =\frac {34}x\\ x= \frac {34} {\cos(27)}=38.1591 \]